if 2y+x+3=0 touches the circle 5X^2 + 5y^2=k


then k= ?

Grade:12

1 Answers

Sanjeev Malik IIT BHU
12 years ago

x2+ y2 =k/5;  centre is (0,0). and radius = root(k/5).

as line touches circle so distance of line from centre equals the radius,

3/root(22+12)= root(k/5)

solving, k= 9